OpenStudy (anonymous):

simplify f(x+h) - f(x) / h f(x)=x^2+x

OpenStudy (anonymous):

ah the difference quotient so beloved by math teachers. only hard part is figuring out what \[f(x+h)\] means. replace x by x+h to get \[f(x+h)=(x+h)^2+(x+h)=x^2+2xh+h^2+x+h\]

OpenStudy (anonymous):

then subtract and divide. everything with out an h in it will disappear from the numerator \[\frac{f(x+h)-f(x)}{h}=\frac{x^2+2xh+h^2+x+h-(x^2+x)}{h}\] \[=frac{x^2+2xh+h^2+x+h-x^2-x}{h}=\frac{2xh+h+h^2}{h}\]

OpenStudy (anonymous):

now factor and cancel to get \[\frac{f(x+h)-f(x)}{h}=\frac{h(2x+1+h)}{h}=2x+1+h\]

OpenStudy (anonymous):

ah great thx

OpenStudy (anonymous):

typo above should be \[=\frac{x^2+2xh+h^2+x+h-x^2-x}{h}=\frac{2xh+h+h^2}{h}\]

OpenStudy (anonymous):

yw